# IBD-PSC: Input-level Backdoor Detection via Parameter-oriented Scaling Consistency

### v2

@@ Appendix D Why Scale the Later Layers?
Our defense relies on building a profile of how the target model behavior changes under progressive modifications to the model. Motivated by a widely accepted hypothesis (e.g., (Tishby & Zaslavsky, 2015; Huang et al., 2022; Jebreel et al., 2023)) that layers situated towards the later stages exert a more direct influence on the ultimate model output, we designed our defense by amplifying the model parameters in stages, starting from the last hidden layer and progressively moving backward through the preceding layers. Here, we examine the alternative of a forward model scaling approach, which scales model parameters starting from the initial layers of the model and progresses forward to the latter layers.
@@ G.1 Backdoor Attacks
  • Training-controlled Backdoor Attacks: The training-controlled attacks include the (8) Dynamic (Nguyen & Tran, 2020), (9) WaNet (Nguyen & Tran, 2021), (10) BPP (Wang et al., 2022b) attacks, and physical backdoor attacks, including (11) PhysicalBA (Li et al., 2021c) and (12) BATT (Xu et al., 2023).
@@ Appendix H Implementation of the Baseline Defenses
Appendix I ROC Curve Comparison with Baseline Defenses
@@
Appendix J Generalizability to Other Model Architectures
Table A7: The performance (AUROC, F1) of the our defense on other model architectures.
@@
Table A8 presents the performance (AUROC and F1 scores) of our IBD-PSC against some other types of backdoor attacks, including clean-label attacks (LC (Turner et al., 2019), NARCISSUS (Zeng et al., 2023)), source-specific attack (TaCT (Tang et al., 2021)), training-controlled attack (BPP (Wang et al., 2022b)), model-controlled attack (SRA (Qi et al., 2022)), and adaptive attack (Adap-Patch (Qi et al., 2023)). The results demonstrate that IBD-PSC consistently outperforms other defense strategies across almost all types of backdoor attacks. It achieves the highest average scores in both AUROC and F1 metrics, marked in bold, underscoring its superior detection capabilities. This comprehensive evaluation affirms the robustness of IBD-PSC as a formidable defense mechanism in the ever-evolving landscape of backdoor attacks in cybersecurity.
@@
Appendix L Settings for the Inference Time Comparison
The inference time is critical for this task (i.e., detecting poisoned testing images) because the detection is usually deployed as the ‘firewall’ for online inference.
In the case of STRIP, TeCo, SCALE-UP, and our defense, defenders utilize the target model’s prediction for defense purposes. This means that both input identification and prediction can be carried out simultaneously.
We calculated the inference time of all defense methods under identical and ideal conditions for evaluating efficiency. For example, we assume that defenders will load all required models and images simultaneously (with more memory requirements compared to the vanilla model inference). This comparison is relatively fair and reasonable since different defenses differ greatly in their mechanisms and requirements. More precisely, before inference, we engage in preparatory steps such as selecting the BN layers to be amplified and preparing the parameter-amplified models. These models are subsequently deployed across different machines, enabling simultaneous processing of input samples. While this approach requires additional storage space to accommodate the various model versions, it considerably accelerates the detection process. For SCALE-UP, we calculate the inference time needed to obtain predictions for multiple augmented images associated with a given input. This is achieved by concurrently feeding all the images into the deployed model as a batch instead of predicting them individually.
Appendix M Ablation Studies
M.1 Ablation Studies on the Threshold $T$
@@
M.2 Ablation Studies on the Hyperparameter $\xi$.
@@
Figure A5: The impact of error rate $\xi$ on defense effectiveness.
(a) BadNets
(b) WaNet
(c) BATT
@@
M.3 Ablation Studies on the Number of Amplified Models $n$
@@
M.4 Robustness to the Target Class
@@
M.5 Ablation Studies on the Size of Local Benign Samples $\mathcal{D}_{r}$
M.6 False Positive Rates on Benign Models
To achieve this objective, we train five benign models using different random seeds. Subsequently, we conduct tests on these models to calculate the false positive rate of our defense, which represents the proportion of benign samples incorrectly identified as backdoor samples. These benign samples are incorrectly rejected by the defense during inference.
@@
Table A9 presents the false positive rates of our defense on different benign models trained on various datasets. We can observe variations in the false positive rate among different models, but overall, it remains relatively low (below 3% on the CIFAR-10 and 1% on the SubImageNet-200 dataset, and around 6% on the GTSRB dataset).
We attribute the higher false positive rate on GTSRB to the relative simplicity of image features in this dataset, making the models more prone to overfitting. Consequently, when amplifying model weights, some benign samples may not decrease their prediction confidence due to overfitting. This overfitting phenomenon on GTSRB has also been reported in the SCALE-UP defense (Guo et al., 2023b). However, in the real world, datasets are often more similar to the ImageNet dataset, characterized by its comprehensive and rich feature information. Our defense performs best on the SubImageNet-200 dataset, achieving an error rate of less than 1%.
In summary, these results indicate that our defense strategy effectively identifies poisoned samples while maintaining a minimal impact on the regular functionality of benign models.
M.7 Is Shrinking as Effective as Amplifying?
According to Theorem 3.1, larger enough feature norms can induce decreasing of confidence in the original predicted class, if the inputs are benign samples (and certain classical assumptions in learning theory are adopted). Poisoned samples, instead, will stay fine. Therefore, by inversely reducing the values of parameters, we expect to observe a degradation in detection performance. The experiments are conducted across a range of reduced magnification factors (0.1 to 0.9) against BadNet, WaNet, and BATT attacks on the CIFAR-10 dataset using the ResNet18 model. The results displayed in Figure A9, clearly indicate a reduction in detection performance, as evidenced by lowered AUROC values and F1 scores approaching zero. This trend remains consistent across the various attack methods examined.
@@
we conduct additional experiments under small poisoning rates (i.e., 0.5% and 1%). Specifically, following the suggestion from the backdoor-toolbox, we remove samples that contain trigger patterns but still cannot be correctly predicted as the target label by attacked DNNs. As shown in Table A10, our defense is still highly effective in these cases, although its performance is slightly lower than that of TeCo, which requires significantly more inference time. In contrast, both STRIP and SCALE-UP failed.
Furthermore, we evaluate the effectiveness of our defense against backdoor attacks with small poisoning rates on the SubImageNet-200 dataset. As demonstrated in Table A11, our defense is still highly effective in defending against attacks with different poisoning rates on the SubImageNet-200 dataset (instead of only effective on CIFAR-10).
Additionally, we compared the performance of our defense with baseline defenses under small poisoning rates. As shown in Table A12, our defense outperforms the baseline defenses in most scenarios across different poisoning rates.
$\rho$ (%) $\rightarrow$ 0.5 1
Defenses $\downarrow$ AUROC TPR FPR AUROC TPR FPR
STRIP 0.403 0.039 0.100 0.421 0.033 0.100
TeCo 1.000 1.000 0.156 1.000 1.000 0.068
SCALE-UP 0.461 0.440 0.389 0.467 0.424 0.345
Ours 0.936 0.791 0.864 0.953 1.000 0.129
@@
Table A13: The attack performance (BA, ASR) with the adaptive attack settings.
@@
Table A14: The effectiveness of the adaptive attack suggested by the reviewer and the performance (AUROC, F1) of IBD-PSC against the adaptive attack on CIFAR-10. We mark the failed cases (where $BA<70\%$) in red, given that the accuracy of models unaffected by backdoor attacks on clean samples is 94.40%.
@@
Design 2. Compared with that ensuring the correct prediction of benign samples under model-parameter amplification in Design 1, we can also design another adaptive attack by decreasing the confidence of predicting poisoned samples to the targeted class by parameter-amplified models. Inspired by label smoothing, we design another adaptive loss term $\mathcal{L}^{\prime}_{\textrm{ada}}$ to decrease the confidence of poisoned samples when model parameter amplification occurs. $\mathcal{L}^{\prime}_{\textrm{ada}}$ is defined as:
@@
where $\hat{y}_{i}$ represents the label-smoothing form of the target class $t$. $\hat{y}_{i}$ is defined as:
@@
Here, $\zeta$ is set to 0.2, specifically chosen to lower the confidence with which poisoned samples are classified into the target class. The term $C$ denotes the total number of classes, $|\mathcal{D}_{p}|$ represents the number of poisoned samples in the training set, and $\boldsymbol{x}_{j}$ denotes a poisoned sample.
We then integrate this adaptive loss term $\mathcal{L}_{\textrm{ada}}$ with the vanilla backdoor loss $\mathcal{L}_{\textrm{bd}}$ to formulate the overall loss function as $\mathcal{L}=\alpha\mathcal{L}_{\textrm{bd}}+(1-\alpha^{\prime})\mathcal{L}_{ada}^{\prime},$ where $\alpha^{\prime}$ is a weighting factor. We evaluate the robustness of our defense under the same settings as those used in our 5.4. As shown in the Table A14, decreasing the confidence of poisoned samples significantly reduces BA and thus the attack can be easily noticed. Furthermore, our defense remains effective even under this new adaptive attack. As analyzed in previous discussions on adaptive attacks, the effectiveness of IBD-SPC largely stems from our adaptive layer selection strategy. This strategy dynamically identifies BN layers for amplification, regardless of whether the model is a vanilla or an adaptively backdoored model, ensuring the robustness of our defense mechanism across various scenarios.
Appendix O How Model Amplification Changes the Latent Representation
@@
Appendix P Comparing Our Defense with Mask-Aware SPC (Pal et al., 2024)
Pal et al. (Pal et al., 2024) introduced a mask-aware scaled prediction consistency (SPC) technique to identify and filter out poisoned samples within the training set. It presents intriguing observations similar to ours regarding SCALE-UP. However, it’s important to clarify that the findings related to SCALE-UP in our paper constitute only a minor component of our research scope. Our study, while similar to the SCALE-UP framework, diverges significantly by exploring the concept from a different angle—parameter scaling. This distinction underlines a substantial difference between our approach and MSPC. Nevertheless, we highlight several critical differences regarding the application scenario of our work, which differs from the training set purification, focused on by MSPC.
Our method requires fewer assumptions about potential adversaries. We explore scenarios where the user employs a third-party model and necessitates a real-time detection of poisoned samples during the inference phase. This detection setup is akin to a firewall and aligns with the setting proposed in SCALE-UP. Importantly, we do not limit our adversaries to employing poison-only attack methods, which is required by MSPC.
Differences in Detection Focus. Our defense is implemented during the inference stage, necessitating the capability for real-time detection. In contrast, training set purification methods are relatively more relaxed regarding detection time since they operate during the data collection phase.
Appendix Q Training set Purification
We also endeavor to apply our defense within the application scenario of identifying and filtering poisoned samples within the training set.
@@
Training set purification aims to filter out potential poisoned samples from a contaminated training set, thereby ensuring that the model trained on the purified dataset is free from backdoors. Many existing studies assume that backdoored models will develop abnormal latent representations for poisoned samples, which are significantly different from those of benign samples, allowing poisoned samples to be identified. Chen et al. (Chen et al., 2018) firstly observe that samples within the target class form two separate clusters in the feature space of the penultimate layer. They employ cluster analysis techniques, such as K-means, to segregate the two clusters. Samples from the smaller cluster are classified as poisoned, based on the assumption that the number of poisoned samples is significantly lower than that of benign samples. Subsequent works generally utilize different cluster analysis methods like singular value decomposition (SVD) (Tran et al., 2018; Hayase et al., 2021), Gram matrix (Ma et al., 2022), K-Nearest-Neighbors (Peri et al., 2020), and feature decomposition (Tang et al., 2021) to detect poisoned samples. Another line of work proposes using differentiating characteristics, such as the faster speed of model fitting (Li et al., 2021b), the presence of high-frequency artifacts (Zeng et al., 2021), and the sensitivity of poisoned samples to transformations (Chen et al., 2022), as methods to identify poiosned samples.
@@
More recently, Huang et al. (Huang et al., 2023) hypothesis that poisoned samples require less input information to be predicted correctly. They introduced the cognitive pattern signature technique, which distills a minimal pattern (given by a mask) for an input sample to retain its original prediction. This technique reveals that poisoned samples typically exhibit a significantly smaller L1 norm of the cognitive pattern than benign samples. Pan et al. (Pan et al., 2023) propose an proactive training set purification method called ASSET, which induces maximum loss difference between poisoned and benign samples by optimizing opposite objectives on the base and poisoned sets. Pal et al. (Pal et al., 2024) present intriguing observations about the limitations of the SCALE-UP (Guo et al., 2023b), based on which they proposed a masked scaled prediction consistency (SPC) technique. This method selectively amplifies specific pixels in input samples, thereby more effectively exposing the prediction invariance of poisoned data to an input scaling factor.
Q.2 Identifying and Filtering Potentially Poisoned Samples within Training set
Following the methodology outlined in references(Pal et al., 2024; Huang et al., 2023; Pan et al., 2023), we first train a model on a potentially compromised training set and subsequently apply our detection method to identify and filter potentially poisoned samples within that training set.
@@
The detection performance is presented in Table A15, which demonstrates the effectiveness of our method in filtering training set samples across various attacks, achieving a 100% TPR and nearly 100% AUROC while maintaining an FPR close to 0%. Note that we reproduce MSPC using its open-source codes with their default settings. However, it performs relatively poorly in defending against WaNet compared to the results reported in its original paper. We speculate that this is probably because we exploited the WaNet with noise mode, whereas MSPC was tested on the vanilla WaNet (as mentioned in their Appendix E). After removing suspected poisoned samples from the training set, we retrain the model on this purified training set to evaluate both its BA and ASR. We conduct experiments on the CIFAR-10 dataset against three representative attacks, and the results are presented in Table A16. As can be seen, the ASR scores of these retrained models are less than 0.5%, thereby rendering these backdoor attacks ineffective.
Appendix R Reproducibility Statement
Appendix S Discussions about the Adopted Data
Appendix T Potential Limitations and Future Directions
@@
Firstly, our defense requires more memory and inference times than the standard model inference without any defense. Specifically, let $M_{s}$ and $M_{d}$ denote the memory (for loading models) required by the standard model inference and by that of our defense, respectively. Let $T_{s}$ and $T_{d}$ denote the inference time required by the standard model inference and by that of our defense. Assuming that we adopt $n$ ($e.g.$, $n=5$) parameter-amplified models for our defense. We have the following equation: $M_{d}\cdot T_{d}=n\times M_{s}\cdot T_{s}.$ Accordingly, the users may need more GPUs to load all/some amplified models simultaneously to ensure efficiency or require more time for prediction by loading those models one by one when the memory is limited. In particular, the storage costs of our defense are similar to those without defense since we can easily obtain amplified models based on the standard one and, therefore, only need to save one model copy ($e.g.$, vanilla model). We will explore how to reduce those costs in our future works.
@@
Fourthly, our work currently focuses only on image classification tasks. We will explore its performance on other modalities (e.g., text and audio) and tasks (e.g., detection and tracking) in our future works.

### v3

@@ Appendix D Why Scale the Later Layers?
Our defense relies on building a profile of how the target model behavior changes under progressive modifications to the model. Motivated by a widely accepted hypothesis (e.g., (Tishby & Zaslavsky, 2015; Huang et al., 2022; Jebreel et al., 2023)) that layers situated towards the later stages exert a more direct influence on the ultimate model output, we designed our defense by amplifying the model parameters in stages, starting from the last hidden layer and progressively moving backward through the preceding layers. Here, we examine the alternative of a forward model scaling approach, which scales model parameters starting from the initial layers of the model and then progressing forward to the latter layers.
@@ G.1 Backdoor Attacks
  • Training-controlled Backdoor Attacks: The training-controlled attacks include the (8) Dynamic (Nguyen & Tran, 2020), (9) WaNet (Nguyen & Tran, 2021), (10) BPP (Wang et al., 2022c) attacks, and physical backdoor attacks, including (11) PhysicalBA (Li et al., 2021c) and (12) BATT (Xu et al., 2023).
@@ Appendix H Implementation of the Baseline Defenses
Appendix I Generalizability to Other Model Architectures
Appendix J ROC Curve Comparison with Baseline Defenses
@@ Appendix I Generalizability to Other Model Architectures
Table A8 presents the performance (AUROC and F1 scores) of our IBD-PSC against some other types of backdoor attacks, including clean-label attacks (LC (Turner et al., 2019), NARCISSUS (Zeng et al., 2023)), source-specific attack (TaCT (Tang et al., 2021)), training-controlled attack (BPP (Wang et al., 2022c)), model-controlled attack (SRA (Qi et al., 2022)), and adaptive attack (Adap-Patch (Qi et al., 2023)). The results demonstrate that IBD-PSC consistently outperforms other defense strategies across almost all types of backdoor attacks. It achieves the highest average scores in both AUROC and F1 metrics, marked in bold, underscoring its superior detection capabilities. This comprehensive evaluation affirms the robustness of IBD-PSC as a formidable defense mechanism in the ever-evolving landscape of backdoor attacks in cybersecurity.
Appendix L Settings for the Inference Time Comparison
The inference time is critical for this task (i.e., detecting poisoned testing images) because the detection is usually deployed as the ‘firewall’ for online inference. In the case of STRIP, TeCo, SCALE-UP, and our defense, defenders utilize the target model’s prediction for defense purposes. This means that both detection and prediction can be carried out simultaneously.
We calculate the inference time of all defense methods under identical and ideal conditions to evaluate efficiency. For example, we assume that defenders will load all required models and images simultaneously, which demands more memory requirements compared to the standard model inference. This comparison is fair and reasonable due to the significant differences in mechanisms and requirements among the various defenses. More precisely, before inference, we engage in preparatory steps such as selecting the BN layers to be amplified and preparing the parameter-amplified models. These models are subsequently deployed across different machines, enabling simultaneous processing of input samples. While this approach requires additional storage space to accommodate the various model versions, it considerably accelerates the detection process. For SCALE-UP, we calculate the inference time needed to obtain predictions for multiple augmented images associated with a given input. This is achieved by concurrently feeding all the images into the deployed model as a batch instead of predicting them individually.
Table A7: The performance (AUROC, F1) of our defense on other model architectures.
@@
(a) BadNets
(b) WaNet
(c) BATT
Figure A5: The impact of error rate $\xi$ on defense effectiveness.
Appendix M Ablation Studies
M.1 Impact of the Threshold $T$
M.2 Impact of the Hyperparameter $\xi$
@@
M.3 Impact of the Number of Amplified Models $n$
@@
M.4 Impact of the Target Class
@@
M.5 Impact of the Size of Local Benign Samples $\mathcal{D}_{r}$
M.6 False Positive Rates on Benign Models
To achieve this objective, we train five benign models using different random seeds. Subsequently, we conduct tests on these models to calculate the false positive rate of our defense, which represents the proportion of benign samples incorrectly identified as backdoor samples. These benign samples are incorrectly rejected by the defense during inference. Table A9 presents the false positive rates of our defense on different benign models trained on various datasets. We can observe variations in the false positive rate among different models, but overall, it remains relatively low (below 3% on the CIFAR-10 and 1% on the SubImageNet-200 dataset, and around 6% on the GTSRB dataset).
We attribute the higher false positive rate on GTSRB to the relative simplicity of image features in this dataset, making the models more prone to overfitting. Consequently, when amplifying model weights, some benign samples may not decrease their prediction confidence due to overfitting. This overfitting phenomenon on GTSRB has also been reported in the SCALE-UP defense (Guo et al., 2023b). However, in the real world, datasets are often more similar to the ImageNet dataset, characterized by its comprehensive and rich feature information. Our defense performs best on SubImageNet-200, achieving an error rate of less than 1%.
@@
M.7 Is Shrinking as Effective as Amplifying?
According to Theorem 3.1, larger enough feature norms can induce a decrease in confidence for the original predicted class, if the inputs are benign samples (and certain classical assumptions in learning theory are adopted). Poisoned samples, instead, will stay fine. Therefore, by inversely reducing the values of parameters, we expect to observe a degradation in detection performance. The experiments are conducted across a range of reduced magnification factors (0.1 to 0.9) against BadNet, WaNet, and BATT attacks on the CIFAR-10 dataset using the ResNet18 model. The results displayed in Figure A9, clearly indicate a reduction in detection performance, as evidenced by lowered AUROC values and F1 scores approaching zero. This trend remains consistent across the various attack methods examined.
@@
We evaluate the robustness of our defense against backdoor attacks with low poisoning rates. Specifically, we focus on three representative attacks (BadNets, WaNet, and BATT) on the CIFAR-10 dataset with poisoning rates ($\rho$) ranging from 0.5% to 10%, and on the SubImageNet-200 dataset with $\rho$ from 2% to 10%, ensuring that most ASRs exceed 80%. The results presented in Table A10 and Table A11 demonstrate that our defense remains effective under low poisoning rates, achieving AUROC and F1 scores well above 0.9, even at a poisoning rate as low as 0.5%. We also compare the performance of our defense with baseline defenses under these conditions. As shown in Table A12, our defense outperforms the baseline defenses in most scenarios across different poisoning rates.
Table A10: The performance of our defense in defending against backdoor attacks with different poisoning rates on CIFAR-10.
Attacks $\rightarrow$ BadNets WaNet BATT
$\rho$downarrow ASR AUROC F1 ASR AUROC F1 ASR AUROC F1
0.5% 1.000 0.955 0.950 0.039 0.936 0.927 0.961 0.913 0.927
1% 1.000 0.950 0.951 0.286 0.953 0.935 0.993 0.948 0.948
2% 1.000 0.999 0.928 0.968 0.966 0.944 0.999 0.999 0.972
4% 1.000 0.998 0.912 0.972 0.977 0.959 1.000 1.000 0.967
6% 1.000 0.999 0.961 0.994 0.978 0.955 1.000 0.998 0.942
8% 1.000 1.000 0.981 0.996 0.983 0.960 1.000 0.985 0.958
10% 1.000 1.000 0.967 0.997 0.984 0.956 1.000 0.999 0.979
@@
We observe that the ASRs (excluding testing samples from the target class) for WaNet are 3.9% and 28.6% at poisoning rates of 0.5% and 1%, respectively. This indicates that the attack nearly fails at a 0.5% poisoning rate. Following the suggestion from the Backdoor-Toolbox, we remove samples containing trigger patterns that still cannot be correctly predicted as the target label by backdoored DNNs. As shown in Table A13, our defense remains highly effective in these cases, although its performance is slightly lower than that of TeCo, which requires significantly more inference time. In contrast, both STRIP and SCALE-UP fail.
Table A13: The comparison of our defense and the baseline defenses in defending against the WaNet attack under poisoning rates of 0.5% and 1% on CIFAR-10.
$\rho$ (%) $\rightarrow$ 0.5 1
Defenses $\downarrow$ AUROC TPR FPR AUROC TPR FPR
STRIP 0.403 0.039 0.100 0.421 0.033 0.100
TeCo 1.000 1.000 0.156 1.000 1.000 0.068
SCALE-UP 0.461 0.440 0.389 0.467 0.424 0.345
Ours 0.936 0.791 0.864 0.953 1.000 0.129
@@
Table A14: The attack performance (BA, ASR) with the adaptive attack settings in “Design 1”.
@@
Table A15: The attack performance (BA, ASR) of the adaptive attack in “Design 2” and the detection performance (AUROC, F1) of IBD-PSC against the adaptive attack on CIFAR-10. We mark the failed cases (where $BA<70\%$) in red, given that the accuracy of models unaffected by backdoor attacks on clean samples is 94.40%.
@@
Table A16: The proportion (%) of misclassified benign samples classified by the model on each category. In our cases, the target label is 0.
Attacks $\downarrow$, Labels $\rightarrow$ 0 1 2 3 4 5 6 7 8 9
BadNets (Original) 14.68 5.71 11.09 22.02 10.77 14.19 6.36 3.92 5.55 5.71
BadNets (Adaptive) 99.57 0.42 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01
BATT (Original) 10.81 5.50 13.18 22.27 10.62 13.18 6.23 5.88 5.78 6.54
BATT (Adaptive) 94.67 0.02 0.30 3.84 0.91 0.09 0.06 0.00 0.02 0.09
Design 2. We can also design another form of adaptive attack to mitigate the impact of parameter amplification. Specifically, we aim to reduce the confidence with which parameter-amplified models predict poisoned samples as belonging to the target class. Inspired by label smoothing, we design an adaptive loss term $\mathcal{L}^{\prime}_{\textrm{ada}}$ to decrease the confidence of poisoned samples when model parameter amplification occurs. The adaptive loss term $\mathcal{L}^{\prime}_{\textrm{ada}}$ is defined as:
@@
where $\hat{y}_{i}$ represents the label-smoothing form of $t$, and $\mathcal{D}_{p}$ denotes the set of poisoned samples. $\hat{y}_{i}$ is defined as:
@@
Here, $\zeta$ is set to 0.2, specifically chosen to reduce the confidence with which poisoned samples are classified into the target class. The term $C$ denotes the total number of classes, $|\mathcal{D}_{p}|$ represents the number of poisoned samples in the training set, and $\boldsymbol{x}_{j}$ denotes a poisoned sample.
We integrate the adaptive loss term $\mathcal{L}^{\prime}_{\textrm{ada}}$ with the vanilla backdoor loss $\mathcal{L}_{\textrm{bd}}$ to formulate the overall loss function as $\mathcal{L}^{\prime}=\alpha\mathcal{L}_{\textrm{bd}}+(1-\alpha^{\prime})\mathcal{L}_{ada}^{\prime}$, where $\alpha^{\prime}$ is a weighting factor. We evaluate the robustness of our defense under the same settings as described in Section 5.4. As shown in Table A15, reducing the confidence of poisoned samples significantly decreases BA, making the attack more noticeable. Moreover, our defense remains effective even against this new adaptive attack. The effectiveness of IBD-SPC largely stems from our adaptive layer selection strategy, which dynamically identifies BN layers for amplification. This approach ensures the robustness of our defense mechanism across various scenarios, whether the model is vanilla or adaptively backdoored.
(a) BadNets
(b) Blend
(c) BATT
(d) Ada-patch
Figure A11: The heatmap of the average prediction confidences of benign samples within the target class predicted by the $n$ scaled models on the CIFAR-10 dataset. The threshold of our defense is set at 0.9; samples with PSC values above this threshold are classified as poisoned. In our cases, the target label is 0.
We claim that the reduction of BA caused by the adaptive attack is mainly because the DNNs link both benign features and trigger features to the target class when using label smoothing, although the connection between trigger features and the target class is stronger. Specifically, attacked models tend to overfit trigger features when high confidence (e.g., 1) is applied on poisoned samples, as seen in vanilla backdoor attacks. However, after label smoothing, the attacked models rely on both trigger and other features, as the task becomes more complicated and harder to fit accurately. Consequently, the adaptive-attacked model is more likely to predict benign samples as the target class, resulting in relatively low benign accuracy. To further verify this, we calculate the distribution of misclassified benign samples. As shown in Table A16, almost all misclassified benign samples are predicted as the target label (i.e., 0) instead of other classes.
We also design an adaptive attack variant that directly classifies poisoned samples to their ground-truth labels under parameter amplification. However, this approach imposes excessive regularization, preventing the learning of backdoors. This phenomenon aligns with our findings in Theorem 3.1. Additionally, we design another adaptive loss term by making parameter-amplified models categorize poisoned images as a different class (rather than the target class). However, even with a very small trade-off hyper-parameter for the adaptive loss term, it significantly decreases the BA of the attacked model by more than $30\%$.
Appendix O The Impact of Parameter Amplification on the Benign Samples from Target class
To better understand the reduction in confidence scores of benign samples, we randomly select 30 benign samples from the target class and generate a heatmap to display the distribution of average logits across $n$ scaled models. As shown in Figure A11, the heatmaps for various attacks highlight a clear clustering phenomenon: scaled models consistently increase prediction confidence for a specific non-original label, thereby reducing confidence scores for the original labels of these benign samples. Moreover, these non-original labels correspond to categories that are more challenging to classify in CIFAR-10 dataset, typically having lower accuracy. For example, the category associated with Logit_3 is “Cat,” which has an accuracy of less than 90%, noticeably below the dataset’s average accuracy of approximately 94%.
@@
Appendix P How Model Amplification Changes the Latent Representation
@@
Appendix Q The Extension to Training Set Purification
@@
Training set purification aims to filter out potentially poisoned samples from a contaminated training set, ensuring that the model trained on the purified dataset is free from backdoors. Many existing studies assume that backdoored models will develop abnormal latent representations for poisoned samples, which are significantly different from those of benign samples, allowing for the identification of poisoned samples. Chen et al. (Chen et al., 2018) firstly observed that samples within the target class form two separate clusters in the feature space of the penultimate layer. They employed cluster analysis techniques, such as K-means, to segregate these clusters. Samples from the smaller cluster are classified as poisoned, based on the assumption that the number of poisoned samples is significantly lower than that of benign samples. Subsequent works generally utilized different cluster analysis methods, such as singular value decomposition (SVD) (Tran et al., 2018; Hayase et al., 2021), Gram matrix (Ma et al., 2022), K-Nearest-Neighbors (Peri et al., 2020), and feature decomposition (Tang et al., 2021), to detect poisoned samples. Another line of research focuses on identifying poisoned samples based on differentiating characteristics, such as the faster speed of model fitting (Li et al., 2021b), the presence of high-frequency artifacts (Zeng et al., 2021), and the sensitivity of poisoned samples to transformations (Chen et al., 2022).
@@
More recently, Huang et al. (Huang et al., 2023) hypothesized that poisoned samples require less input information to be predicted correctly. They introduced the cognitive pattern signature technique, which distills a minimal pattern (given by a mask) for an input sample to retain its original prediction. This technique reveals that poisoned samples typically exhibit a significantly smaller L1 norm in the cognitive pattern compared to benign samples. Pan et al. (Pan et al., 2023) proposed a proactive training set purification method called ASSET, which maximizes the loss difference between poisoned and benign samples by optimizing opposite objectives on the base and poisoned sets. Pal et al. (Pal et al., 2024) presented intriguing observations regarding the limitations of SCALE-UP (Guo et al., 2023b), leading to the proposal of the masked scaled prediction consistency (MSPC) technique. This method selectively amplifies specific pixels in input samples, thereby more effectively exposing the prediction invariance of poisoned data under an input scaling factor.
Q.2 Comparing Our IBD-PSC with MSPC
MSPC (Pal et al., 2024) presents observations that are similar to ours regarding SCALE-UP. However, it is important to clarify that our findings on SCALE-UP constitute only a minor component of our research. While our study shares some similarities with the SCALE-UP framework, it significantly diverges by exploring parameter scaling, highlighting a substantial difference from MSPC. Additionally, our work focuses on different application scenarios compared to MSPC, which is primarily concerned with training set purification.
Our Method Requires Fewer Assumptions about Potential Adversaries. We explore scenarios where the users employ third-party models and need real-time detection of poisoned samples during the inference phase, similar to a firewall. This setup aligns with the framework proposed in SCALE-UP. Notably, we do not limit our adversaries to using poison-only attack methods, which is required by MSPC.
Differences in Detection Focus. Our defense operates during the inference stage, requiring the capability of real-time detection. In contrast, MSPC is less constrained in detection time as it operates during the data collection phase.
Q.3 Identifying and Filtering Potentially Poisoned Samples within Training set
@@
Following the methodology in references (Pal et al., 2024; Huang et al., 2023; Pan et al., 2023), we first train a model on a potentially compromised training set and then apply our detection method to identify and filter potentially poisoned samples within that dataset. The detection performance, presented in Table A17, demonstrates the effectiveness of our method in filtering malicious training samples across various attacks, achieving a 100% TPR and nearly 100% AUROC while maintaining an FPR close to 0%. Note that we reproduce MSPC using its open-source codes with default settings. However, it performs relatively poorly in defending against WaNet compared to the results reported in its original paper. We speculate this is because we test WaNet in noise mode, whereas MSPC is tested on the vanilla WaNet (as mentioned in their Appendix E). After removing suspected poisoned samples from the training set, we retrain the model on this purified training set to evaluate both its BA and ASR. We conduct experiments on the CIFAR-10 dataset against three representative attacks, and the results, presented in Table A18, show that the ASR scores of these retrained models are less than 0.5%, rendering these backdoor attacks ineffective.
Appendix R Potential Limitations and Future Directions
@@
Firstly, our defense requires more memory and inference times than the standard model inference without any defense. Specifically, let $M_{s}$ and $M_{d}$ denote the memory (for loading models) required by the standard model inference and by that of our defense, respectively. Let $T_{s}$ and $T_{d}$ denote the inference time required by the standard model inference and by that of our defense. Assuming that we adopt $n$ (e.g., $n=5$) parameter-amplified models for our defense. We have the following equation: $M_{d}\cdot T_{d}=n\times M_{s}\cdot T_{s}.$ Accordingly, the users may need more GPUs to load all/some amplified models simultaneously to ensure efficiency or require more time for prediction by loading those models one by one when the memory is limited. In particular, the storage costs of our defense are similar to those without defense since we can easily obtain amplified models based on the standard one and, therefore, only need to save one model copy (e.g., vanilla model). We will explore how to reduce those costs in our future work.
@@
Fourthly, our work currently focuses only on image classification tasks. We will explore its performance on other modalities (e.g., text and audio) and tasks (e.g., detection and tracking) in our future work.
Appendix S Reproducibility Statement
Appendix T Discussions about the Adopted Data